Three charges lie on X-axis: 1x10 С at x 3x10*C 1 cm, 2 x 10.8 С at x-2 cm, and 3. at x -3 cm. What is the potential energy of this system?
Answer #1

here,

q1 = 1 * 10^-8 C is at x1 = 1 cm = 0.01 m

q2 = 2 * 10^-8 C is at x2 = 2 cm = 0.02 m

q3 = 3 * 10^-8 C is at x3 = 3 cm = 0.03 m

the potential energy of this system , PE = K * q1 *q2/(x2 - x1) + K * q2 * q3/(x3 - x2) + K * q1 * q3 /( x3 - x1)

PE = 9 * 10^9 * 10^-8 * 10^-8 * ( 1*2/(0.02 - 0.01) + 2 * 3 /( 0.03 - 0.02) + 1 * 3 /(0.03 - 0.01)) J

PE = 8.55 * 10^-4 J

the potential energy of this system is 8.55 * 10^-4 J